OpenStudy (anonymous):

A company makes similar cylindrical containers in two sizes, as shown in the table below. Containers Radius(in Cm) Volume(in Cubic Cm) Type 1 4 ?? Type 2 7 1231.66 What is the volume, in cubic centimeters, of Type 1 containers?

OpenStudy (ganpat):

v cylinder = pi * r2 * h so for type 2 = 1231.66 = pi * r2 * h therefore h = (1231.66 )/ (pi * r2) FOr Type 1 = V = pi * 4 sq * 1231.66 / (pi * 7 sq) So V = 402.174 cm cube..